OpenStudy (anonymous):

If the slope of a line is 5/3 and it passes through points (6,4) and (x,8) what is the value of x... The answer is 42/5 ( I only know that cuz my summer assignment answer sheet told me) but I have to show detailed work and I don't know what to do!! Please help me!!!

OpenStudy (nubeer):

do you know how to find the equation of line?

OpenStudy (anonymous):

I think so.... Is that the x-x1 or something... Don't make fun of me if its not, it's just a guess:)

OpenStudy (nubeer):

yup something like that.. you can find the equation of line by this formula y-y1= m (x-x1) u have the slope.. and point (6,4)

OpenStudy (anonymous):

Wait, and stay with me here, does that help. So I think this is where I'm supposed to do this: 5-3 ------- 6-4 Is this right?

OpenStudy (nubeer):

Ummm no O.o you have the slope.. m = 5/3 and point (6,4) out these values in the equation i wrote in the previous comment.

OpenStudy (anonymous):

Oh sorry... Okay did that

OpenStudy (nubeer):

so what is the equation you got?

OpenStudy (anonymous):

Is it 6-4= 5-3? If not I'm not being stupid on purpose

OpenStudy (nubeer):

y-4 = 5/3 (x -6) now try solving it

OpenStudy (anonymous):

Can someone please help me? I'm desperate!!!,:)